A precipitate could form when you mix Ca+2and F-1? What is the value of Qsp for the precipitate solubility when 0.275 L of 0.30M Ca2+ is mixed with 0.204 L of 0.060 M F-1?

**Remeber Volumes are additive since it is a mixture
Answer #1

Answer

i) A precipitate could form

ii) Qsp = 1.13×10-4

Explanation

Solubility equillibrium of CaF2 is

CaF2(s) <------> Ca2+(aq) + 2F-(aq)

Ksp = [Ca2+] [F-]2 = 3.9 ×10-11

​​​​​​Concentration after mixing

[Ca2+] = 0.30M / (0.479L/0.275L ) = 0.1722M

[F-] = 0.060M/(0.479L/0.204L) = 0.0256M

Qsp = [Ca2+][F-]2

Qap = 0.17222 × (0.0256)2

Qsp = 1.13×10-4

Qsp > Ksp

Therefore,

Precipitate will form